OpenStudy (anonymous):

can someone teach me how to do this Solve x + 3y = 9 3x − 3y = −13

OpenStudy (anonymous):

Hi Welcome to Openstudy! Are you familiar with the substitution and elimination methods?

OpenStudy (anonymous):

kinda but its been awhile ahaha

OpenStudy (anonymous):

Hey can I get a medal for the last question that I answered @pascuzzzo I worked really hard on that

OpenStudy (anonymous):

Substitution is when you re-arrange one of the equation in terms a variable and on the 2nd equation, you replace that variable by the new equation that you get. Elimination method is when you manipulate the given equation in order to eliminate one variable to solve for the other variable. Now, from your equations, which one do you want to do? Elimination method or substitution? any method will work. just pick one and we will work it out step by step ^_^

OpenStudy (anonymous):

@jackmullen55 I figured it out myself tbh and substitution

OpenStudy (anonymous):

oh thanks... still I worked really hard mate

OpenStudy (anonymous):

okay :) First, let's label each equation: x + 3y = 9 -> eqn 1 3x − 3y = −13 -> eqn 2 2nd, let's rearrange equation 1 in terms of x. Rearrange means move 3y on the right side. Can you do that?

OpenStudy (anonymous):

3x=-13+3y?

OpenStudy (anonymous):

it is easier to do it with eqn 1: x +3 y = 9 (try changing this instead of the eqn 2) because eqn 2 will be in fraction, which will be pretty messy.

OpenStudy (anonymous):

x=9+3y

OpenStudy (anonymous):

quite right, but it should be : x = 9-3y because you have to subtract both sides by 3y to move it to the right side. Next, we have eqn 2: \(\sf 3 \color{red}{x} - 3y = -13\) Replace the x variable in eqn 2 from the new equation in eqn1 \(\sf \color{red}{x = 9-3y}\) so it will be: \(\sf 3(\color{red}{ 9-3y}) - 3y = -13\) Now, simplify it and solve for y.

OpenStudy (anonymous):

oh wow I feel dumb I got it now thank you

OpenStudy (anonymous):

:) ok goodluck!